# Large-eddy simulations of heat transfer from iced surfaces

### v1

@@ II.3 Treatment of solid thermal transients
Figure 2: Root-mean-square roughness height ($k_{rms}$), 2N-point-average ($N=5$) peak-to-trough roughness height ($k_{z}$), and skewness ($Sk$) as a function of the streamwise coordinate for the three geometries.
III Computational setup
@@
The computational setup corresponds to the experiments of Ref. 23, conducted in the Baylor University Subsonic Wind Tunnel. In this study, the geometries labeled 113012.05, 112912.02, and 113012.04 will be considered. The geometries correspond to the unwrapped suction section of iced NACA 0012 airfoils created in the Icing Research Tunnel at NASA Glenn Research Center. They result from exposing the airfoils at a freestream temperature of $-2.37\degree$C to supercooled large droplets (SLD) conditions for 43, 55, and 75 seconds, respectively. The iced geometries were laser scanned and unwrapped using the methodology of Ref. 31. Hereafter, the geometries, depicted in Figure 1 and colored by surface elevation, will be referred by their exposure time as $s43$, $s55$, and $s75$. The $x$, $y$, and $z$ coordinates correspond to the streamwise, wall-normal, and spanwise directions, respectively.
[…]
@@
The rough surfaces were 3D printed in ABS plastic, placed over a neoprene layer and mounted over a Plexiglass support. In between the Plexiglass and the neoprene, five Mylar plates provided heating to the surface, with an approximate heat flux of $q_{\text{gen}}\approx$ 500 W/m2. Thermocouples were installed at the Mylar plate, at the bottom of the plexiglas layer, and in the free-stream flow. They were used to estimate the heating lost through the Plexiglass base ($q_{\text{HL}}$). A schematic of the experimental setup, showing the four layers present in the experiment and the heat fluxes, is presented in Figure 3. In the heat transfer experiments, the geometry was scaled by a factor of 10 and the testing velocity was reduced proportionally to maintain the Reynolds number achieved during the accretion process.
@@
Figure 3: Schematic of the experimental setup, showing the four layers present in the experiment and the heat fluxes (adapted from Ref. 23)
@@
The computational domain includes the totality of the plate, which is $L_{x}=37.5$ inches long and $L_{z}=8$ inches wide. The boundary layer is not tripped, resulting in a laminar boundary layer that undergoes transition caused by the beginning of the rough region. The transition location ($X_{t}$), which coincides approximately with the beginning of roughness, obtained from skin friction measurements performed on the experiments, is $X_{t}/L_{x}=0.4$, $0.3$, and $0.41$ for the $s43$, $s55$, and $s75$ geometries, respectively. On the fluid domain, the inlet is set with a mean free-stream velocity of $U_{\infty}=6.67$ m/s, and for the outlet a non-reflecting characteristic boundary condition was used [33]. Slip boundary conditions are used at the top and the side walls. At the plate, wall models are used to calculate the shear stress. For $x\leq X_{t}$ a laminar wall model is employed, while for $x>X_{t}$ an algebraic equilibrium wall model is used. The temperature at the surface results from the solution of the heat equation in the solid domain. The free-stream temperature is set based on the experiments to $T_{\infty}=293.15\degree$K. The solid domain considers only the neoprene and plastic layers. The material properties for neoprene are specified as $\rho_{n}=1230$ kg/m3, $c_{p,n}=2200$ J/(kg·K), and $\lambda_{s,n}=0.175$ W/(m·K). For ABS plastic, the properties are set to $\rho_{p}=1070$ kg/m3, $c_{p,n}=1990$ J/(kg·K), and $\lambda_{s,n}=0.162$ W/(m·K). A sensitivity analysis involving a 20% variation in the material properties of neoprene and ABS plastic was conducted, showing only minor variations in surface temperatures (within 5% of the experimental uncertainty). At the bottom boundary, constant heat fluxes for each heating plate are set based on heat generated in the plates minus the heat lost through the Plexiglass base $(q_{\text{gen}}-q_{\text{HL}})$. The heating on each individual panel is calculated based on the reported current and voltage on the individual heaters as $q_{\text{gen}}=EV/A_{h}$, where $E$ is the current through heater, $V$ is the voltage across heater, and $A_{h}$ is the area of the heater. Lateral walls are set using adiabatic boundary conditions. At the interface with the fluid domain, heat fluxes are imposed based on the solution of the fluid domain and the radiative fluxes.
Geometry Grid $N_{\mathrm{cv}}$ [Mcv] $k_{\mathrm{rms}}/\Delta_{\min}$ $k_{z}/\Delta_{\min}$ $\delta/\Delta_{\min}$
(fluid–solid) (max–mean) (max–mean) ($0.1L_{x}$$0.4L_{x}$$0.7L_{x}$)
s43 Coarse 1.5–0.2 1.1–0.6 5.6–3.7 1.9–8–17
Medium 6.3–2.0 2.2–1.1 11–7.4 3.9–15–34
Fine 25–15 4.5–2.3 22–15 7.7–30–67
s55 Coarse 1.6–0.2 1.9–0.9 7.9–5.7 1.9–13–22
Medium 6.6–1.9 3.7–1.8 16–11 3.9–27–44
Fine 27–16 7.5–3.6 32–23 7.7–54–88
s75 Coarse 1.7–0.2 2.6–1.3 11–7.3 1.9–10–24
Medium 6.9–2.2 5.2–2.6 21–15 3.9–19–49
Fine 28–17 11–5.2 42–29 7.7–38–97
[…]
@@
The body-fitted computational mesh based on Voronoi diagrams is constructed using the Lloyd algorithm [34]. Three meshes of sequential refinements levels are used in the simulations. The details of the meshes for the three geometries are presented in Table 1. Each refinement level is achieved via halving the characteristic cell size. The meshes are refined at the surface and progressively coarsened away from it by reducing one refinement level every 15 layers. In addition to the wall-normal clustering, using isotropic grid cells ensures adequate resolution of the roughness elements in the streamwise and spanwise directions. The fine mesh resolution resulted in an average of 2.3–5.2 grid elements per $k_{\text{rms}}$ and 15–29 elements per $k_{z}$. Figure 4 presents images of the fine mesh over two planes, highlighting the resolution near the rough surface (for the fluid mesh) and inside a roughness element (for the solid mesh). These resolutions allow us to capture the flow around roughness elements as well as temperature variations within them. As shown in Table 1, the resolution within the laminar region of the boundary layer ($\delta/\Delta_{\text{min}}$ at $0.1L_{x}$) is insufficient to accurately resolve the thin thermal laminar boundary layer. However, as demonstrated in Section IV, this limitation affects only a small portion of the domain and does not compromise the accuracy of the downstream results. Increasing the resolution in this region would significantly increase the computational cost without notable benefit.
IV Validation
@@
To validate the model, temperature distributions over the plate surface are compared with the experimental data. Surface temperature data is used to validate the model because it is the only variable related to heat transfer directly measured in the experiments. Other quantities, such as heat transfer coefficient or $St$, were derived from the temperature measurements. Figure 5 presents spanwise-averaged surface temperature profiles over a central band of thickness $\Delta z/L_{z}=1/2$ for the three geometries, comparing numerical predictions with IR camera measurements. Spanwise averaging is restricted to the central half of the plate to minimize the influence of boundary effects observed in the experiments. To maintain clarity, results for the three mesh resolutions are shown only for geometry $s43$, which exhibits the smallest roughness height, and results for the finest mesh are shown for $s55$ and $s75$. Adequate resolution for $s43$ case is expected to ensure sufficient resolution for the other geometries, which feature larger roughness elements. Results from the medium and fine meshes for $s43$ are closely aligned, indicating mesh convergence in surface temperature predictions. For the finer meshes, the computed temperatures generally fall within the experimental uncertainty (±$\pm 1.4\degree$K) across most of the domain. Local maxima and minima in the temperature profile are accurately captured in the roughness region, highlighting the model’s effectiveness in predicting heat transfer over rough surfaces.
@@
All three geometries exhibit discrepancies with the experimental data in the laminar region ($x/L_{x}\leq 0.2$), likely due to a combination of insufficient resolution in the early portion of the boundary layer and the presence of small pressure gradients induced by the leading edge of the mounting structure, which is not included in the numerical model. Since the primary focus of this study is on heat transfer in the rough region, these differences are considered acceptable. Resolving the thin laminar boundary layer in detail would require a significantly finer mesh than that employed here, with substantial computational cost increase [35]. Moreover, the observed discrepancies in the upstream laminar region do not affect the results in the downstream region of interest. In addition, case $s75$ presents some discrepancies with the experimental data near $x/L_{x}=0.4$, where the numerical model overpredicts the wall temperature peak at the onset of the roughness region. These differences diminish quickly within the initial portion of the rough surface. These results show that a mesh resolution of approximately $k_{\text{rms}}/\Delta\approx 2$ is enough to capture heat transfer and temperature variations over ice roughness. These results are in line with resolution requirements to predict aerodynamic loads on iced airfoils [22]. Overall, the model shows strong agreement with experimental observations, indicating its high accuracy in representing heat transfer features.
@@
Figure 5: Spanwise-averaged surface temperature profiles over a band of thickness $\Delta z/L_{z}=1/2$ located at the plate centerline obtained with the coarse, medium, and fine meshes, as well as with the IR camera for geometry $s43$ and with fine mesh, for $s55$ and $s75$. $\langle$.$\rangle$ . ⟩ operator indicates spanwise-averaging. The gray are indicates the uncertainty of the measurements with the IR camera ($1.4\degree$K). Dashed line represents spanwise averaged surface elevation.
@@
[…]
@@ V Roughness characterizaton
where $B$ depends on the value of $k_{s}^{+}$. In the fully rough regime, $B$ approaches a constant value that is a function of the roughness geometry. The superscript ‘+’ denotes variables normalized using inner (viscous) scaling. The wall-normal distance is corrected by the zero-plane displacement, which is estimated as the center of drag on the wall [39]. Alternatively, velocity profiles can be expressed using the roughness function $\Delta U^{+}$, which quantifies the downward shift of the log law relative to a smooth wall at matched wall-normal distance $y^{+}$:
@@
The existence of a logarithmic layer requires that $y>y_{k}$ and $y\ll\delta$ (or $y_{k}\ll\delta$), where $y_{k}$ is the height of the roughness sublayer, defined as the region where the flow is directly influenced by the local roughness topology, and is typically estimated as $y_{k}\approx 3k_{s}$ [40]. The existence of outer-layer similarity has been historically associated with small roughness-to-boundary-layer-height ratios. Jiménez [41] suggested that outer-layer similarity holds for $k/\delta<1/40$, although subsequent studies have demonstrated its validity for $k_{z}/\delta<1/19$ [40], and even up to $k/\delta\approx 1/7$ [18].
Figure 8: $k_{z}/\delta$ ratio as a function of the streamwise coordinate for cases $s43$, $s55$ and $s75$
Figure 8 shows the variation of $k_{z}/\delta$ along the plate for the three geometries. Values of $k_{z}/\delta$ reach approximately $0.6-0.8$ in regions where the roughness height is largest, and then decrease downstream as the boundary layer thickens. These values far exceed the thresholds typically associated with the existence of outer-layer similarity or even a logarithmic layer. Additional velocity deficit profiles (presented in Appendix B) confirm that similarity is only recovered in the outer wake region in the final portion of the plate in cases $s43$ and $s55$. The high $k_{z}/\delta$ ratios indicate that the flow over these surfaces is more representative of an obstacle regime rather than a canonical rough-wall regime. In this regime, the influence of roughness elements extends across the entire boundary layer, and parameters such as $\Delta U^{+}$ and $k_{s}^{+}$ no longer provide a meaningful description of the velocity profiles.
As noted at the beginning of this section, $k_{s}$ is widely used in icing simulations. For that reason, even though this parameter may not be physically meaningful in the present context, we still compute an equivalent sand-grain roughness, that is, a value of $k_{s}$ that would yield the same wall shear stress as the actual rough surface. Since a logarithmic layer is absent in many of the velocity profiles, Equation (10) cannot be used to estimate $k_{s}^{+}$ directly. However, Equation (11) can be rearranged to express the roughness function $\Delta U^{+}$ in terms of the difference between friction coefficients for smooth and rough walls at matched $\delta^{+}$:
@@
To compute $\Delta U^{+}$ using these models, spanwise-averaged roughness statistics are employed as inputs. A filter with a length of $L_{x}/20$ is applied to remove high-frequency oscillations in roughness statistics resulting from the limted domain span. All three models tend to overpredict the values of $k_{s}^{+}$, which is expected given that they were originally calibrated for significantly lower values of $k/\delta$ than those present in the current simulations. The models proposed by Bornhoft et al. [43] and De Marchis et al. [44] capture the overall trend of the computed $k_{s}^{+}$ values reasonably well. In particular, they reproduce the sharp rise at the beginning of the rough region, the location of the peak at $(x-X_{t})/L_{x}=0.2$, and the gradual decay toward negligible values by $(x-X_{t})/L_{x}=0.53$. However, both models fail to predict the secondary peak observed at $(x-X_{t})/L_{x}=0.4$. In contrast, the model from Flack and Schultz [42] captures this secondary peak but substantially overestimates $k_{s}^{+}$ for $(x-X_{t})/L_{x}>0.2$. Figure 9 (right) shows the values of $C_{f}$ obtained directly from simulations (accounting for both skin friction and form drag) alongside those estimated indirectly from the $k_{s}^{+}$ values predicted by the previously discussed models. As expected, the overestimation of $k_{s}^{+}$ leads directly to an overestimation of $C_{f}$ on the rough portion of the plate. The largest discrepancy was observed using Equation (14), with an average error of approximately 138%. The models by Bornhoft et al. [43] and De Marchis et al. [44] performed comparatively better, yielding mean estimation errors in $C_{f}$ of 61% and 43%, respectively. These results highlight that under the present conditions, predicting surface roughness via available $k_{s}^{+}$ models can result in significant inaccuracies in estimating both $k_{s}^{+}$ and consequently wall-shear stress. It is important to emphasize that this limitation is not inherent to the models themselves; rather, these models were specifically developed for rough-wall conditions where outer-layer similarity holds. In icing scenarios, roughness heights can grow rapidly and reach magnitudes comparable to the boundary layer thickness, thus violating the outer-layer similarity hypothesis, and significantly reducing the accuracy of conventional roughness parametrizations. Hence the approach of using smooth surfaces superimposed with sand-grain roughness may not be very appropriate for simulating icing conditions.

### v2

@@ II.4 Treatment of solid thermal transients
Figure 2: Root-mean-square roughness height ($k_{rms}$), 2N-point-average ($N=5$) peak-to-trough roughness height ($k_{z}$), skewness ($Sk$), and effective slope ($ES$) as a function of the streamwise coordinate for the three geometries.
III Experimental data
@@
The experimental data used to validate the model correspond to the experiments described in Ref. 25, which were conducted in the Baylor University Subsonic Wind Tunnel. In this study, the geometries labeled 113012.05, 112912.02, and 113012.04 will be considered. The geometries correspond to the unwrapped suction section of iced NACA 0012 airfoils created in the Icing Research Tunnel at NASA Glenn Research Center. They result from exposing the airfoils at a freestream temperature of $2.37\degreeC$ to supercooled large droplets (SLD) conditions for 43, 55, and 75 seconds, respectively. The iced geometries were laser scanned and unwrapped using the methodology of Ref. 35. The three unwrapped geometries are depicted in Figure 1 and colored by surface elevation. Throughout this article, $x$, $y$, and $z$ coordinates represent to the streamwise, wall-normal, and spanwise directions, respectively. Hereafter, the geometries will be referred by their exposure time as $s43$, $s55$, and $s75$.
@@
The rough surfaces were 3D printed in ABS plastic, placed over a neoprene layer and mounted over a Plexiglass support. In between the Plexiglass and the neoprene, five Mylar plates provided heating to the surface, with an approximate heat flux of $q_{\text{gen}}\approx$ 500 W/m2. Thermocouples were installed at the Mylar plate, at the bottom of the plexiglas layer, and in the free-stream flow. They were used to estimate the heating lost through the Plexiglass base ($q_{\text{HL}}$). A schematic of the experimental setup, showing the four layers present in the experiment and the heat fluxes, is presented in Figure 3. In the heat transfer experiments, the geometry was scaled by a factor of 10, and the testing velocity was reduced proportionally to maintain the Reynolds number achieved during the accretion process. The three plates are $L_{x}=37.5$ inches long and $L_{z}=8$ inches wide. The free-stream velocity was set to $U_{\infty}=6.67$ m/s. The Reynolds and Mach numbers of the experiment are $Re_{Lx}\approx 4.2\times 10^{5}$ and $Ma\approx 0.02$. The boundary layer is not tripped, resulting in a laminar boundary layer that undergoes transition caused by the beginning of the rough region. The transition location ($X_{t}$), which coincides approximately with the beginning of roughness, obtained from skin friction measurements, is $X_{t}/L_{x}=0.4$, $0.3$, and $0.41$ for the $s43$, $s55$, and $s75$ geometries, respectively.
@@
Figure 3: Schematic of the experimental setup, showing the four layers present in the experiment and the heat fluxes (adapted from Ref. 25)
Each surface initially exhibits a smooth region followed by the onset of roughness at different streamwise locations. These differences arise from the stochastic nature of ice formation, as each surface represents a distinct realization under identical experimental conditions. Figure 2 shows the streamwise evolution of spanwise averaged root-mean-square roughness height ($k_{\text{rms}}$), the spanwise averaged 2NN-point-average ($N=5$) peak-to-trough roughness height ($k_{z}$), spanwise averaged skewness ($Sk$), and spanwise averaged effective slope ($ES$) for the three geometries. For definitions of these parameters, the reader is referred to Ref. 22. Small frequency variations of these parameters, resulting from the limited width of the domain, are removed by applying a filter of length $L_{x}/20$. The roughness height reaches its maximum near the beginning of the rough region and progressively decreases downstream. The maximum roughness height is located at $X_{k,\max}/L_{x}=0.44,0.37$ and 0.5 for $s43$, $s55$, and $s75$ geometries, respectively, with peaks of $k_{\text{rms}}$ and $k_{z}$ coinciding. In $s43$ and $s55$, the surface transitions back to an almost flat plate towards the downstream end. In contrast, roughness is still present at the end of $s75$. For all geometries, the faster decrease of $k_{\text{rms}}$ compared to $k_{z}$, along with the associated increase in skewness, indicates that roughness elements become more scattered towards the downstream region. Such spatial variability in roughness height distribution is characteristic of ice accretion processes [36].
We have selected this dataset to validate our model because it provides high-resolution measurements of temperature distributions at various accretion times. This enables us to evaluate the model at different stages and to assess the effects of changing roughness heights. However, it is important to acknowledge certain differences between these experiments and actual icing conditions. First, in real icing, heat is released at the surface primarily due to the latent heat of solidification, whereas in the experiments considered here, heat is supplied from below the surface. This may affect the role of conduction on the heat-transfer process. Second, the experiments are conducted at a free-stream temperature of $293.15~\mathrm{K}$ ($20^{\circ}$C), while icing typically occurs at much lower ambient temperatures. Finally, the thermal conductivity of ABS plastic used in the experiments is nearly an order of magnitude lower than that of ice. Although both are low-conductivity materials, this difference can lead to varying temperature gradients and alter the significance of conductive heat transfer within the solid.
Despite these limitations, we consider this dataset to be highly valuable for studying heat transfer over surfaces with ice-like roughness. It enables us to validate convective heat transfer calculations under different accretion stages and to investigate the role of the solid’s conduction on the convective heat transfer for low-conductivity materials.
IV Computational setup
@@
The computational setup replicates the experimental setup described on the previous section. A schematic of the computational domain, including boundary conditions, is presented in Figure 4. The fluid domain includes the totality of the plate, while the solid domain includes the plastic and neoprene layers. On the fluid domain, the inlet is set with the mean free-stream velocity, and for the outlet a non-reflecting characteristic boundary condition is used [37]. Slip boundary conditions are used at the top and the side walls. At the plate, wall models are used to calculate the shear stress. For $x\leq X_{t}$ a Blasius boundary condition is employed, while for $x>X_{t}$ an algebraic equilibrium wall model is used. The temperature at the surface results from the solution of the heat equation in the solid domain. The free-stream temperature is set based on the experiments to $T_{\infty}=293.15$ K. The solid domain considers only the neoprene and plastic layers. The material properties for neoprene are specified as $\rho_{n}=1230$ kg/m3, $c_{p,n}=2200$ J/(kg·K), and $\lambda_{s,n}=0.175$ W/(m·K). For ABS plastic, the properties are set to $\rho_{p}=1070$ kg/m3, $c_{p,p}=1990$ J/(kg·K), and $\lambda_{s,p}=0.162$ W/(m·K). The ratio of solid to fluid conductivity is $\lambda_{s,n}/\lambda=7$ and $\lambda_{s,p}/\lambda=6.5$. A sensitivity analysis involving a 20% variation in the material properties of neoprene and ABS plastic was conducted, showing only minor variations in surface temperatures (within 5% of the experimental uncertainty). At the bottom boundary, constant heat fluxes for each heating plate are set based on heat generated in the plates minus the heat lost through the Plexiglass base $(q_{\text{gen}}-q_{\text{HL}})$. The heating on each individual panel is calculated based on the reported current and voltage on the individual heaters as $q_{\text{gen}}=EV/A_{h}$, where $E$ is the current through heater, $V$ is the voltage across heater, and $A_{h}$ is the area of the heater. Lateral walls are set using adiabatic boundary conditions. At the interface with the fluid domain, heat fluxes are imposed based on the solution of the fluid domain and the radiative fluxes.
Geometry Grid $N_{\mathrm{cv}}$ [Mcv] $k_{\mathrm{rms}}/\Delta_{\min}$ $k_{z}/\Delta_{\min}$ $\delta_{99}/\Delta_{\min}$ $y^{+}$
[fluid, solid] [max, mean] [max, mean] $(0.1,\,0.4,\,0.7)\,L_{x}$] [max, mean]
s43 Coarse 1.5, 0.2 1.1, 0.6 5.6, 3.7 1.9, 8, 17 68, 20
Medium 6.3, 2.0 2.2, 1.1 11, 7.4 3.9, 15, 34 34, 10
Fine 25, 15 4.5, 2.3 22, 15 7.7, 30, 67 17, 5
s55 Coarse 1.6, 0.2 1.9, 0.9 7.9, 5.7 1.9, 13, 22 68, 20
Medium 6.6, 1.9 3.7, 1.8 16, 11 3.9, 27, 44 34, 10
Fine 27, 16 7.5, 3.6 32, 23 7.7, 54, 88 17, 5
s75 Coarse 1.7, 0.2 2.6, 1.3 11, 7.3 1.9, 10, 24 70, 22
Medium 6.9, 2.2 5.2, 2.6 21, 15 3.9, 19, 49 35, 11
Fine 28, 17 11, 5.2 42, 29 7.7, 38, 97 17, 6
Table 1: Grid refinement details: cell counts $N_{\mathrm{cv}}$ in the fluid and solid domain; ratio of mean and maximum root‐mean‐square roughness height ($k_{\mathrm{rms}}$) and two‐point‐average peak‐to‐trough roughness height ($k_{z}$) to minimum grid length scale ($\Delta_{\min}$); and ratio of boundary layer height ($\delta_{99}$) to $\Delta_{\min}$ at $x/L_{x}=0.1,\,0.4,\,0.7$.
Figure 4: Schematic of the computational domain. Red text indicates the fluid domain boundary conditions, and blue text indicates the solid domain boundary conditions. Vertical dimensions are not to scale.
@@
The body-fitted computational mesh based on Voronoi diagrams is constructed using the Lloyd algorithm [38]. Three meshes of sequential refinements levels are used in the simulations. The details of the meshes for the three geometries are presented in Table 1. Each refinement level is achieved via halving the characteristic cell size. The meshes are refined at the surface and progressively coarsened away from it by reducing one refinement level every 15 layers. In addition to the wall-normal clustering, using isotropic grid cells ensures adequate resolution of the roughness elements in the streamwise and spanwise directions. The fine mesh resulted in an average of $k_{\text{rms}}/\Delta_{\min}=2.3-5.2$ grid elements per $k_{\text{rms}}$ and 15–29 elements per $k_{z}$, where $\Delta_{\min}$ represent the minimum cell size (located at the wall). Figure 5 presents images of the fine mesh over two planes, highlighting the resolution near the rough surface (for the fluid mesh) and inside a roughness element (for the solid mesh). These resolutions allow us to capture the flow around roughness elements as well as temperature variations within them. As shown in Table 1, the resolution within the laminar region of the boundary layer ($\delta_{99}/\Delta_{\text{min}}$ at $0.1L_{x}$) is insufficient to accurately resolve the thin thermal laminar boundary layer. However, as demonstrated in Section V, this limitation affects only a small portion of the domain and does not compromise the accuracy of the downstream results. Increasing the resolution in this region would significantly increase the computational cost without notable benefit.
V Validation
@@
To validate the model, temperature distributions over the plate surface are compared with the experimental data. Surface temperature data is used to validate the model because it is the only variable related to heat transfer directly measured in the experiments. Other quantities, such as heat transfer coefficient or $St$, were derived from the temperature measurements. Figure 6 presents spanwise-averaged surface temperature profiles over a central band of thickness $\Delta z/L_{z}=1/2$ for the three geometries, comparing numerical predictions with IR camera measurements. Spanwise averaging is restricted to the central half of the plate to minimize the influence of boundary effects observed in the experiments. To maintain clarity, results for the three mesh resolutions are shown only for geometry $s43$, which exhibits the smallest roughness height, and results for the finest mesh are shown for $s55$ and $s75$. Adequate resolution for $s43$ case is expected to ensure sufficient resolution for the other geometries, which feature larger roughness elements. Results from the medium and fine meshes for $s43$ are closely aligned, indicating mesh convergence in surface temperature predictions. For the finer meshes, the computed temperatures generally fall within the experimental uncertainty ($\pm 1.4$ K) across most of the domain. Local maxima and minima in the temperature profile are accurately captured in the roughness region, highlighting the model’s effectiveness in predicting heat transfer over rough surfaces.
@@
All three geometries exhibit discrepancies with the experimental data in the laminar region ($x/L_{x}\leq 0.2$), likely due to a combination of insufficient resolution in the early portion of the boundary layer and the presence of small pressure gradients induced by the leading edge of the mounting structure, which is not included in the numerical model. Since the primary focus of this study is on heat transfer in the rough region, these differences are considered acceptable. Resolving the thin laminar boundary layer in detail would require a significantly finer mesh than that employed here, with substantial computational cost increase [39]. Moreover, the observed discrepancies in the upstream laminar region do not affect the results in the downstream region of interest. In addition, case $s75$ presents some discrepancies with the experimental data near $x/L_{x}=0.4$, where the numerical model overpredicts the wall temperature peak at the onset of the roughness region. These differences diminish quickly within the initial portion of the rough surface. These results show that a mesh resolution of approximately $k_{\text{rms}}/\Delta_{\min}\approx 2$ is enough to capture heat transfer and temperature variations over ice roughness. These results are in line with resolution requirements to predict aerodynamic loads on iced airfoils [24]. Overall, the model shows strong agreement with experimental observations, indicating its high accuracy in representing heat transfer features.
@@
Figure 6: Spanwise-averaged surface temperature profiles over a band of thickness $\Delta z/L_{z}=1/2$ located at the plate centerline obtained with the coarse, medium, and fine meshes, as well as with the IR camera for geometry $s43$ and with fine mesh, for $s55$ and $s75$. $\langle.\rangle$ operator indicates spanwise-averaging. The gray are indicates the uncertainty of the measurements with the IR camera ($\pm 1.4$ K). Dashed line represents spanwise averaged surface elevation.
@@
Figure 8: Velocity profiles at $x/L_{x}=0.29,\;0.49,$ and 0.68 (left, center, right, respectively) for $s55$ geometry obtained with the numerical model and from the experiments.
@@ VI Roughness characterizaton
where $B$ depends on the value of $k_{s}^{+}$. In the fully rough regime, $B$ approaches a constant value that is a function of the roughness geometry. $\kappa$ is the von Kármán constant, and the superscript ‘+’ denotes variables normalized using inner (viscous) scaling. The wall-normal distance is corrected by the zero-plane displacement, which is estimated as the center of drag on the wall [43]. Alternatively, velocity profiles can be expressed using the roughness function $\Delta U^{+}$, which quantifies the downward shift of the log law relative to a smooth wall at matched wall-normal distance $y^{+}$:
@@
The existence of a logarithmic layer requires that $y>y_{k}$ and $y\ll\delta$ (or $y_{k}\ll\delta$), where $y_{k}$ is the height of the roughness sublayer, defined as the region where the flow is directly influenced by the local roughness topology, and is typically estimated as $y_{k}\approx 3k_{s}$ [44]. The existence of outer-layer similarity has been historically associated with small roughness-to-boundary-layer-height ratios. Jiménez [45] suggested that outer-layer similarity holds for $k/\delta<1/40$, although subsequent studies have demonstrated its validity for $k/\delta<1/19$ [44], and even up to $k/\delta\approx 1/7$ [22].
Figure 9: $k_{z}/\delta_{99}$ ratio as a function of the streamwise coordinate for cases $s43$, $s55$ and $s75$
Figure 9 shows the variation of $k_{z}/\delta_{99}$ along the plate for the three geometries. Values of $k_{z}/\delta_{99}$ reach approximately 0.60.80.6-0.8 in regions where the roughness height is largest, and then decrease downstream as the boundary layer thickens. These values far exceed the thresholds typically associated with the existence of outer-layer similarity or even a logarithmic layer. Additional velocity deficit profiles (presented in Appendix B) confirm that outer-layer similarity is not recovered within the length of the domain. The high $k_{z}/\delta_{99}$ ratios indicate that the flow over these surfaces is more representative of an obstacle regime rather than a canonical rough-wall regime. In this regime, the influence of roughness elements extends across the entire boundary layer, and parameters such as $\Delta U^{+}$ and $k_{s}^{+}$ no longer provide a meaningful description of the velocity profiles.
As noted at the beginning of this section, $k_{s}$ is widely employed in icing simulations, despite the fact that $k/\delta$ often exceeds the established limits for the existence of outer-layer similarity. For this reason, even though this parameter may not be physically meaningful in the present context, we still compute an equivalent sand-grain roughness, that is, a value of $k_{s}$ that would yield the same wall shear stress as the actual rough surface. Since a logarithmic layer is absent in many of the velocity profiles, Equation (14) cannot be used to estimate $k_{s}^{+}$ directly. However, Equation (15) can be rearranged to express the roughness function $\Delta U^{+}$ in terms of the difference between friction coefficients for smooth and rough walls at matched $\delta^{+}$:
@@
To compute $\Delta U^{+}$ using these models, spanwise-averaged roughness statistics are employed as inputs. A filter with a length of $L_{x}/20$ is applied to remove high-frequency oscillations in roughness statistics resulting from the limited domain span. All three models tend to overpredict the values of $k_{s}^{+}$, which is expected given that they were originally calibrated for significantly lower values of $k/\delta$ than those present in the current simulations. The models proposed by Bornhoft et al. [47] and De Marchis et al. [48] capture the overall trend of the computed $k_{s}^{+}$ values reasonably well. In particular, they reproduce the sharp rise at the beginning of the rough region, the location of the peak at $(x-X_{t})/L_{x}=0.2$, and the gradual decay toward negligible values by $(x-X_{t})/L_{x}=0.53$. However, both models fail to predict the secondary peak observed at $(x-X_{t})/L_{x}=0.4$. In contrast, the model from Flack and Schultz [46] captures this secondary peak but substantially overestimates $k_{s}^{+}$ for $(x-X_{t})/L_{x}>0.2$. Figure 10 (right) shows the values of $C_{f}$ obtained directly from simulations (accounting for both skin friction and form drag) alongside those estimated indirectly from the $k_{s}^{+}$ values predicted by the previously discussed models. As expected, the overestimation of $k_{s}^{+}$ leads directly to an overestimation of $C_{f}$ on the rough portion of the plate. The largest discrepancy was observed using Equation (18), with an average error of approximately 138%. The models by Bornhoft et al. [47] and De Marchis et al. [48] performed comparatively better, yielding mean estimation errors in $C_{f}$ of 61% and 43%, respectively. However, in the final portion of the plate, where the ratio $k_{z}/\delta_{99}$ is smaller, these two models yield lower mean errors of 26% and 23%, respectively. These results highlight that under the present conditions, predicting surface roughness via available $k_{s}^{+}$ models can result in significant inaccuracies in estimating both $k_{s}^{+}$ and consequently wall-shear stress. It is important to emphasize that this limitation is not inherent to the models themselves; rather, these models were specifically developed for rough-wall conditions where outer-layer similarity holds. Furthermore, models to estimate $k_{s}^{+}$ were developed for fully turbulent regimes, not for transitional regimes such as those encountered in the initial portion of the rough region. In icing scenarios, roughness heights can grow rapidly and reach magnitudes comparable to the boundary layer thickness, thus violating the outer-layer similarity hypothesis and significantly reducing the accuracy of conventional roughness parametrizations. Hence the approach of using smooth surfaces superimposed with sand-grain roughness may not be very appropriate for simulating icing conditions.